OpenStudy (anonymous):

-1/2x =-3/4 solve using the multiplication principle

OpenStudy (anonymous):

Multiply both sides by -2. You'll have your answer.

OpenStudy (hoblos):

-1/2x =-3/4 multiply both sides by 2 =>x=3/2

OpenStudy (anonymous):

so do multiply ny 2 or -2 ?

OpenStudy (hoblos):

sorry -2

OpenStudy (anonymous):

thanks hoblos and moorsy!